# Large Language Model Enhanced Knowledge Representation Learning: A Survey

### v3

@@ Abstract
Knowledge Representation Learning (KRL) is crucial for enabling applications of symbolic knowledge from Knowledge Graphs (KGs) to downstream tasks by projecting knowledge facts into vector spaces. Despite their effectiveness in modeling KG structural information, KRL methods are suffering from the sparseness of KGs. The rise of Large Language Models (LLMs) built on the Transformer architecture present promising opportunities for enhancing KRL by incorporating textual information to address information sparsity in KGs. LLM-enhanced KRL methods, including three key approaches, encoder-based methods that leverage detailed contextual information, encoder-decoder-based methods that utilize a unified seq2seq model for comprehensive encoding and decoding, and decoder-based methods that utilize extensive knowledge from large corpora, has significantly advanced the effectiveness and generalization of KRL in addressing a wide range of downstream tasks. This work provides a broad overview of downstream tasks while simultaneously identifying emerging research directions in these evolving domains.
@@ 1 Introduction
Triples in Knowledge Graphs (KGs) are an effective form of structured knowledge representation, consisting of an entity, a relation, and an object, where the object can be either another entity or a literal. This structured fact is crucial for various downstream tasks such as link prediction [1] (e.g., predicting ”Albert Einstein” is associated with ”Physics” through the relation ”field of study”), triple classification [2] (e.g., verifying if ”Paris is the capital of France” is true), and relation classification [3] (e.g., classifying ”wrote” as the relation connecting ”J.K. Rowling” and ”Harry Potter”). For better application of symbolic knowledge from KGs to these tasks, Knowledge Representation Learning (KRL) approaches have been developed. Notable methods like TransE [4], RESCAL [5], and R-GCN [6] focus on embedding information (entities, relations, etc.) into low-dimension vector spaces.
@@
Despite being effective in modelling KG structural information, these methods suffer from the sparseness of KGs. Specifically, some KRL models are trained to preserve the inherent structure of KGs, favoring entities that are highly connected. According to the research, it is widely observed that the degrees of entities in KGs approximately follow the power-law distribution, resulting in a long tail populated with massive unpopular entities of low degrees [7]. Consequently, due to information sparsity, the performance of KRL methods tends to degrade when processing long-tail entities. In short, the sparseness of KGs significantly impacts the effectiveness of representing entities with low degrees is an existing challenge.
@@
To tackle the above challenges, a promising solution is to enhance KRL approaches with large language models. Large Language Models (LLMs), built on the Transformer architecture [8], have gained significant popularity in the NLP field due to their remarkable performance. By pre-trained with vast text corpora, LLMs demonstrate profound content comprehension and rich real-world knowledge, which can be utilized to address the information sparsity in KGs by incorporating textual information such as entity descriptions, offering potential opportunities for better knowledge representations.
@@
Nowadays, the potential of LLM-enhanced KRL methods has attracted increasing attention from academia to industry. In the stages of LLM development, these models exhibit an increase in parameters and real-world knowledge, progress from good to strong understanding capabilities, and transition from task-specific to more generalized outputs. This evolution can be illustrated through the development of three Transformer architectures. Accordingly, LLM-enhanced KRL methods can be categorized into three types based on these architectures. In encoder-based methods, the text context is fully leveraged to capture detailed contextual information for knowledge representation. In encoder-decoder-based methods, a seq2seq model is utilized to intuitively and simply meet all requirements through efficient encoding and decoding. In decoder-based methods, extensive knowledge from large corpora is fully empowered for downstream tasks.
@@
There are few existing surveys exploring the enhancement of KRL by LLM. With respect to KRL, Ge et al. [9] provide a comprehensive overview of the two main branches in KRL: distance-based and semantic matching-based approaches, with few references to LLM enhancement. Biswas et al. [10] give an overview of the advancements in KRL by incorporating features such as semantics, multi-modal, temporal, and multilingual aspects, while rarely address LLM empowerment. Cao et al. [11] classify and analyze KRL models from the perspective of mathematical spaces, specifically through algebraic, geometric, and analytical perspective, excluding considerations related to LLM augmentation. During this period, Pan et al. [12] mainly focus on the integration of LLMs and KGs from the perspective of general frameworks, KG-enhanced LLMs, LLM-augmented KGs, and Synergized LLMs + KGs. They demonstrate how to leverage their respective advantages in a complementary manner, with a minor portion dedicated to LLM-enhanced KRL.
@@ 2.1 Translation Model
Since TransE can only handle one-to-one relations, many extensions have been developed to address one-to-many, many-to-one, and many-to-many relations. For example, Wang et al. proposed TransH [15], in which a relation as a hyperplane. TransR [14] introduces a relation-specific space, and TransD [13] constructs a dynamic projection matrix for each entity and relation, simultaneously considering the diversity of entities and relations.
@@ 3 LLM-Enhanced KRL Methods
Table 1 presents a comparison of various methods utilized for KRL with LLMs, emphasizing their input types, advantages, and challenges. Each method adopts a unique approach to encoding and processing triples or related information, catering to specific tasks and scenarios within the broader scope of KRL. Encoder-based approaches like triple-based representation treats the triple as a single unit, providing a holistic representation but struggling with unseen triples. Translation-based representation pairs the head and relation for improved reasoning, although it may be less expressive. Independent representation handles head, relation, and tail separately, supporting modularity and zero-shot learning but lacking integration across elements. Similarly, encoder-decoder-based structure-based representation methods leverages triple sequences for efficient structural representation, while textual fine-tuning simplifies adaptation by focusing on textual triples but is limited by the generative power of the underlying model. Decoder-based approaches like description generation, which enhances representation in low-resource settings via prompts but is heavily dependent on prompt quality, and prompt engineering, which harnesses the pre-trained knowledge of LLMs but at high computational costs. Lastly, structural fine-tuning combines structural and textual embeddings to balance textual and structural knowledge, though it requires sophisticated integration.
@@ 3.1.1 Triple-based Representation
Figure 1: An overview of triple-based representation method.
@@
where $\text{Tok}{hi}$, $\text{Tok}{ri}$, and $\text{Tok}_{ti}$ represent the tokens of the head entity, relation, and tail entity descriptions, respectively. The model computes the final hidden vector of the unique [CLS] token, which serves as the aggregate sequence representation. The plausibility score of a triple $\tau=(h,r,t)$ is then computed as follows:
@@
PKGC [33] is a pretrained language model (PLM) based model for knowledge graph completion that utilizes prompt engineering to enhance knowledge utilization. This model adapts PLM for the task by converting knowledge triples into natural language prompts, facilitating the effective use of PLM’ latent knowledge.
@@ 3.1.2 Translation-based Representation
Figure 2: An overview of translation encoding method.
@@
Figure 3: An overview of independent representation method.
@@ 3.2.1 Structure-based Representation
Structure-based representation leverages the structure of the triples, feeding them into the encoder as sequences. which allows the model to capture both the syntactic and relational structure of the triples. By integrating structural information with natural language input, this method enhances the ability to represent complex relations. Fig. 4 demonstrates the GenKGC [38] architecture that exemplifies structure-based representation.
Figure 4: An overview of structure-based representation method.
@@ 3.2.2 Textual Fine-Tuning
Textual fine-tuning methods adapt pre-trained encoder-decoder models using natural language descriptions of triples. This approach tailors the model’s generative capabilities to specific knowledge representation tasks, offerring significant advantages in terms of training and inference efficiency. An illustration of a typical textual fine-tuning model KGT5 [34] is presented in Fig. 5.
Figure 5: An overview of fine-tuning method.
@@ 3.3.1 Description Generation
Description generation methods enhance the representation of low-resource entities by creating descriptive text. This approach supplements the textual information, improving the performance of previous methods and ensuring a richer, more complete entity representation. Fig. 6 illustrates a typical method contextualization distillation (CD) [46] used in description generate method.
Figure 6: An overview of description generate method.
@@ 3.3.2 Prompt Engineering
Figure 7: An overview of prompt engineering method.
@@ 3.3.3 Structural Fine-Tuning
Figure 8: An overview of structural fine-tuning method.
@@ 4 Experiments and Evaluations
This section systematically reviews existing experiments and evaluations to analyze the performance and effectiveness of various LLM-enhanced KRL methods discussed previously. By consolidating results from various studies, we aim to present a structured overview of their potential and constraints.
@@ 4.3.5 Link Prediction
As displayed in Table 11, UMLS is a domain-focused, smaller graph where structural baselines already score high Hits@10 (98–99%). LP-BERT, by leveraging textual descriptions, further reduces mean rank to 1.18 and achieves Hits@10 of 100%. In contrast, Wikidata5M is considerably more diverse: DistMult here only reaches an MRR of 25.3%, while LLM-enhanced SimKGC and CSPromp-KG push the MRR to 35.8% and 38.0%, respectively. These improvements underscore how large language models can enrich entity representations, particularly under high relational heterogeneity.
@@ 5 Future Directions
Recent advances in LLMs mark a significant inflection point in KRL research, presenting exciting opportunities for future exploration on effectively integrating the advantages of LLMs into KRL. We propose the following six recommendations:
@@ 5.1 Dynamic and Multimodal Knowledge Representation
As knowledge evolves far beyond static triples, future KRL approaches must accommodate temporal or event-based representations that can handle the fluid nature of real-world data. Modeling sequences of events and capturing causal dependencies enables a more fine-grained understanding of how entities and relations form, merge, or dissolve over time. Such temporal awareness can reveal nuanced patterns that remain invisible under static assumptions, thereby enhancing both accuracy and timeliness of knowledge graphs.

### v5

@@ Abstract
Knowledge Representation Learning (KRL) is crucial for enabling applications of symbolic knowledge from Knowledge Graphs (KGs) to downstream tasks by projecting knowledge facts into vector spaces. Despite their effectiveness in modeling KG structural information, KRL methods are suffering from the sparseness of KGs. The rise of Large Language Models (LLMs) built on the Transformer architecture presents promising opportunities for enhancing KRL by incorporating textual information to address information sparsity in KGs. LLM-enhanced KRL methods, including three key approaches, encoder-based methods that leverage detailed contextual information, encoder-decoder-based methods that utilize a unified Seq2Seq model for comprehensive encoding and decoding, and decoder-based methods that utilize extensive knowledge from large corpora, have significantly advanced the effectiveness and generalization of KRL in addressing a wide range of downstream tasks. This work provides a broad overview of downstream tasks while simultaneously identifying emerging research directions in these evolving domains.
@@ 1 Introduction
Triples in Knowledge Graphs (KGs) are an effective form of structured knowledge representation, consisting of an entity, a relation, and an object, where the object can be either another entity or a literal. This structured fact is crucial for various downstream tasks such as link prediction [1] (e.g., predicting ”Albert Einstein” is associated with ”Physics” through the relation ”field of study”), triple classification [2] (e.g., verifying if ”Paris is the capital of France” is true), and relation classification [3] (e.g., classifying ”wrote” as the relation connecting ”J.K. Rowling” and ”Harry Potter”). To better leverage the symbolic knowledge contained in KGs for these downstream tasks, various Knowledge Representation Learning (KRL) approaches have been developed. Notable methods like TransE [4], RESCAL [5], and R-GCN [6] focus on embedding information (entities, relations, etc.) into low-dimensional vector spaces.
@@
Despite being effective in modelling KG structural information, these methods suffer from the sparseness of KGs. Specifically, some KRL models are trained to preserve the inherent structure of KGs, favoring entities that are highly connected. According to the research, it is widely observed that the degrees of entities in KGs approximately follow the power-law distribution, resulting in a long tail populated with massive unpopular entities of low degrees [7]. Consequently, due to information sparsity, the performance of KRL methods tends to degrade when processing long-tail entities. In short, the sparseness of KGs significantly impacts the effectiveness of of representing low-degree entities, posing an existing challenge.
@@
To tackle the above challenges, a promising solution is to enhance KRL approaches with large language models. Large Language Models (LLMs), built on the Transformer architecture [8], have gained significant popularity in the NLP field due to their remarkable performance. By being pre-trained on vast text corpora, LLMs demonstrate profound content comprehension and rich real-world knowledge, which can be utilized to address the information sparsity in KGs by incorporating textual information such as entity descriptions, offering potential opportunities for better knowledge representations.
@@
Nowadays, the potential of LLM-enhanced KRL methods has attracted increasing attention across both academia and industry. In the stages of LLM development, these models exhibit an increase in parameters and real-world knowledge, progress from good to strong understanding capabilities, and transition from task-specific to more generalized outputs. This evolution can be illustrated through the development of three Transformer architectures. Accordingly, LLM-enhanced KRL methods can be categorized into three types based on these architectures. In encoder-based methods, the textual context is fully leveraged to capture detailed contextual information for knowledge representation. In encoder-decoder-based methods, a seq2seq model is utilized to intuitively and simply meet all requirements through efficient encoding and decoding. In decoder-based methods, extensive knowledge from large corpora is fully harnessed for downstream tasks.
@@
Few existing surveys have explored the enhancement of KRL using LLM. With respect to KRL, Ge et al. [9] provide a comprehensive overview of the two main branches in KRL: distance-based and semantic matching-based approaches, with few references to LLM enhancement. Biswas et al. [10] give an overview of the advancements in KRL by incorporating features such as semantics, multi-modal, temporal, and multilingual aspects, while rarely addressing LLM empowerment. Cao et al. [11] classify and analyze KRL models from the perspectives of mathematical spaces, specifically through algebraic, geometric, and analytical perspective, excluding considerations related to LLM augmentation. During this period, Pan et al. [12] mainly focus on the integration of LLMs and KGs from the perspective of general frameworks, KG-enhanced LLMs, LLM-augmented KGs, and Synergized LLMs + KGs. They demonstrate how to leverage their respective advantages in a complementary manner, with a minor portion dedicated to LLM-enhanced KRL.
@@ 2.1 Translation Model
Since TransE can only handle one-to-one relations, many extensions have been developed to address one-to-many, many-to-one, and many-to-many relations. For example, Wang et al. proposed TransH [15], in which each relation is modeled as a hyperplane. TransR [14] introduces a relation-specific space, and TransD [13] constructs a dynamic projection matrix for each entity and relation, simultaneously considering the diversity of entities and relations.
@@ 3 LLM-Enhanced KRL Methods
Table 1 presents a comparison of various methods utilized for KRL with LLMs, emphasizing their input types, advantages, and challenges. Each method adopts a unique approach to encoding and processing triples or related information, catering to specific tasks and scenarios within the broader scope of KRL. Encoder-based approaches like triple-based representation treats the triple as a single unit, providing a holistic representation but struggling with unseen triples. Translation-based representation pairs the head and relation for improved reasoning, although it may be less expressive. Independent representation handles head, relation, and tail separately, supporting modularity and zero-shot learning but lacking integration across elements. Similarly, encoder-decoder-based structure-based representation methods leverage triple sequences for efficient structural representation, while textual fine-tuning simplifies adaptation by focusing on textual triples but is limited by the generative power of the underlying model. Decoder-based approaches like description generation, which enhances representation in low-resource settings via prompts but is heavily dependent on prompt quality, and prompt engineering, which harnesses the pre-trained knowledge of LLMs but at high computational costs. Lastly, structural fine-tuning combines structural and textual embeddings to balance textual and structural knowledge, though it requires sophisticated integration.
@@ 3.1.1 Triple-based Representation
Figure 1: An overview of triple-based representation methods.
@@
where $\text{Tok}_{hi}$, $\text{Tok}_{ri}$, and $\text{Tok}_{ti}$ represent the tokens of the head entity, relation, and tail entity descriptions, respectively. The model computes the final hidden vector of the unique [CLS] token, which serves as the aggregate sequence representation. The plausibility score of a triple $\tau=(h,r,t)$ is then computed as follows:
@@
PKGC [33] is a pre-trained language model (PLM) based model for knowledge graph completion that utilizes prompt engineering to enhance knowledge utilization. This model adapts PLM for the task by converting knowledge triples into natural language prompts, facilitating the effective use of latent knowledge of the PLM.
@@ 3.1.2 Translation-based Representation
Figure 2: An overview of translation encoding methods.
@@
Figure 3: An overview of independent representation methods.
@@ 3.2.1 Structure-based Representation
Structure-based representation leverages the structure of the triples by feeding them into the encoder as sequences. which allows the model to capture both the syntactic and relational structure of the triples. By integrating structural information with natural language input, this method enhances the ability to represent complex relations. Fig. 4 demonstrates the GenKGC [38] architecture that exemplifies structure-based representation.
Figure 4: An overview of structure-based representation methods.
@@ 3.2.2 Textual Fine-Tuning
Textual fine-tuning methods adapt pre-trained encoder-decoder models using natural language descriptions of triples. This approach tailors the model’s generative capabilities to specific knowledge representation tasks, offering significant advantages in terms of training and inference efficiency. An illustration of a typical textual fine-tuning model KGT5 [34] is presented in Fig. 5.
Figure 5: An overview of fine-tuning methods.
@@ 3.3.1 Description Generation
Description generation methods enhance the representation of low-resource entities by creating descriptive text. This approach supplements the textual information, improving the performance of previous methods and ensuring a richer, more complete entity representation. Fig. 6 illustrates a typical contextualization distillation (CD) [46] method used in description generation method.
Figure 6: An overview of description generation methods.
@@ 3.3.2 Prompt Engineering
Figure 7: An overview of prompt engineering methods.
@@ 3.3.3 Structural Fine-Tuning
Figure 8: An overview of structural fine-tuning methods.
@@ 4 Experiments and Evaluations
This section systematically reviews existing experiments and evaluations to analyze the performance and effectiveness of various LLM-enhanced KRL methods discussed previously. By consolidating results from various studies, we aim to present a structured overview of their potential and constraints. The best and second-best results in tables are shown in bold and underlined, respectively.
@@ 4.3.5 Link Prediction
As listed in Table 11, UMLS is a domain-focused, smaller graph where structural baselines already score high Hits@10 (98–99%). LP-BERT, by leveraging textual descriptions, further reduces mean rank to 1.18 and achieves Hits@10 of 100%. In contrast, Wikidata5M is considerably more diverse: DistMult here only reaches an MRR of 25.3%, while LLM-enhanced SimKGC and CSPromp-KG push the MRR to 35.8% and 38.0%, respectively. These improvements underscore how large language models can enrich entity representations, particularly under high relational heterogeneity.
@@ 5 Future Directions
Recent advances in LLMs mark a significant inflection point in KRL research, presenting exciting opportunities for future exploration of effectively integrating the advantages of LLMs into KRL. We propose the following six recommendations:
@@ 5.1 Dynamic and Multimodal Knowledge Representation
As knowledge evolves far beyond static triples, future KRL approaches must accommodate temporal or event-based representations that can handle the dynamic and evolving characteristics inherent in real-world data. Modeling sequences of events and capturing causal dependencies enables a more fine-grained understanding of how entities and relations form, merge, or dissolve over time. Such temporal awareness can reveal nuanced patterns that remain invisible under static assumptions, thereby enhancing both accuracy and timeliness of knowledge graphs.
@@ 6 Conclusion
[…]
Acknowledgements This work is supported by the National Natural Science Foundation of China (62472311, U23B2057, 62176185).